OpenStudy (anonymous):

If an 10 kg object has a potential energy of 125 J, what velocity would it be right before it hit the ground?

OpenStudy (anonymous):

You might try your luck on physicsforum.

OpenStudy (anonymous):

google it.

OpenStudy (anonymous):

mv^2/2 = 125 J -> v^2 = 250/ 10 = 25 => v = 5